OpenStudy (anonymous):

(4a-3)=a+13/a+1 (4a-3)=a+13/a+1 @Mathematics

OpenStudy (agreene):

\[4a-3=\frac{a+13}{a+1}\] yes? Do you have any thoughts for an answer?

OpenStudy (anonymous):

-1,1?

OpenStudy (agreene):

close, but not quite there :X

OpenStudy (agreene):

multiply both sides by a+1 \[(4a-3)(a+1)=a+13\] expand using FOIL \[4a^2+a-3=a+13\] subtract a and 13 from both sides \[4a^2-16=0\] add 16 to both sides \[4a^2=16\] divide by 4 \[a^2=\frac{16}{4}\] \[a^2=4\] \[a=\pm\sqrt{4}\] \[a=\pm 2\] a= -2;2

OpenStudy (anonymous):

i have no clue where im going wrong

OpenStudy (anonymous):

oh the adding 16 to both sides

OpenStudy (agreene):

well, I can't see your work--but id say look carefully since ive given all of mine on these and try and figure it out :D

OpenStudy (anonymous):

i subtracted 4-3